# FreeAudio: Training-Free Timing Planning for Controllable Long-Form Text-to-Audio Generation

### v1

@@ 3.2 Problem Formulation
We present a training-free method to adapt a pretrained DiT-based T2A model for long-form generation with timing control. We start with a base model, denoted as $f$, which takes a text prompt $y^{c}$ (i.e., overall audio caption) and a target duration $M$ as input to generate an audio waveform of $M$ seconds, where $M\in(0,M_{\text{max}})$. In practice, $M_{\text{max}}$ is typically set to 10 seconds for most existing T2A models. To enable timing control, our method incorporates an additional input: timing prompts $y^{t}=\{(p_{i}^{c},s_{i},e_{i})\}_{i=1}^{n}$. Here, each element $(p_{i}^{c},s_{i},e_{i})$ specifies that the audio content described by the fine-grained text prompt $p_{i}^{c}$ should occur within the time interval $[s_{i},e_{i}]$.
Specifically, for the timing control capability, FreeAudio employs a two-stage planning-and-generation approach: 1) Planning: Leveraging LLM’s capabilities, FreeAudio first generates a detailed timing plan $y^{p}=\{(p_{j}^{c},t_{j},t_{j+1})\}_{j=1}^{k}$, using the overall text prompt $y^{c}$ and the timing prompts $y^{t}$. This plan $y^{p}$ divides the total $M$-second duration into $k$ sequential, non-overlapping windows $\{[t_{j},t_{j+1}]\}_{j=1}^{k}$ with $0=t_{1}<\cdots<t_{k+1}=M$, each associated with a recaptioned prompt $p_{j}^{c}$. 2) Generation: Based on the resulting timing plan, FreeAudio employs the Decoupling & Aggregating Attention Control to manipulate the attention outputs across the time axis, aligning each prompt $p_{j}^{c}$ with its corresponding time window $[t_{j},t_{j+1}]$.
Furthermore, this planning-and-generation schema facilitates long-form audio generation, extending beyond the base model’s maximum duration. Instead of directly modeling the entire target duration $M^{\prime}>M_{\text{max}}$, which exceeds the base model’s capacity, our approach first leverages the generated timing plan $y^{p}$ to partition the task into a sequence of manageable segments, each typically having a duration less than or equal to $M_{\text{max}}$. For smoother transitions between these segments, FreeAudio introduces the Contextual Latent Composition applied at each diffusion sampling step. Additionally, FreeAudio employs the proposed Reference Guidance to facilitate the maintenance of global consistency across the full audio sequence of duration $M\textquoteright$ ’.
@@ 3.3 LLM Planning
As shown in Figure 1 (left), the LLM planning stage takes the text prompt $y^{c}$ and timing prompts $y^{t}$ to generate a detailed timing plan $y^{p}$. This plan consists of a sequence of non-overlapping time windows $[t_{j},t_{j+1}]$, each associated with recaptioned prompts $p_{j}^{c}$. Specifically, we structure this planning stage using chain-of-thought (CoT) [38, 43] reasoning in two steps: timing-to-window planning to establish the time windows, followed by text recaption to assign a detailed prompt to each window.
@@
The timing-to-window planning stage is necessary for two primary reasons: 1) Temporal Overlaps: Input intervals $[s_{i},e_{i}]$ in the timing prompts $y^{t}$ can overlap. Merging the text encoder’s outputs for separate overlapping prompts (e.g., $p_{x}^{c}$ and $p_{y}^{c}$) poses challenges for accurate generation guidance. In comparison, establishing non-overlapping windows shifts the fusion to the text input stage (e.g., concatenating prompts such as $p_{x}^{c}$ and $p_{y}^{c}$), which better leverages the pretrained text encoder’s capability for interpreting unified textual descriptions of combined events. 2) Temporal Gaps: The input timing prompts $y^{t}$ might not cover the entire duration, leaving gaps where no audio event is specified. By leveraging the LLM’s planning ability, this stage infers and assigns appropriate audio events for the uncovered periods within the generated timing plan $y^{p}$. This ensures temporal coherence and semantic consistency in the audio narrative, informed by the context of the planned timing windows and the overall text prompt $y^{c}$.
Specifically, the timing-to-window planning process involves several steps: First, we define the window boundaries by collecting all unique start ($s_{i}$) and end ($e_{i}$) timestamps from $y^{t}$, inserting $t_{1}=0$ and $t_{k+1}=M$, and sorting them to obtain $0=t_{1}<t_{2}<\dots<t_{k+1}=M$. This defines the non-overlapping windows $[t_{j},t_{j+1}]$. Second, we initialize a preliminary description for the window-level events by aggregating all fine-grained text prompts $p_{i}^{c}$ whose original intervals $[s_{i},e_{i}]$ encompass the current window $[t_{j},t_{j+1}]$. Third, leveraging the overall text prompt $y^{c}$ and the current window-level event assignments, the LLM performs a planning step to systematically assign the remaining audio events from $y^{c}$ and distribute them into appropriate window slots, prioritizing the filling of uncovered intervals, thereby finalizing the overall event plan.
@@ 3.4 Timing-Controlled Generation
A naive approach to timing control might independently generate each $(t_{j+1}-t_{j})$-second audio segment using only its associated prompt $p_{j}^{c}$ from the timing plan $y^{p}$ [42]. However, simply concatenating these segments often results in poor inter-window consistency and lacks global guidance from the overall prompt $y^{c}$. To address these limitations, we construct a batch inference pipeline that jointly leverages both $y^{c}$ and $y^{p}$. Specifically, we form a batch of size $1+k$ (one base latent and $k$ sub-latents) for the base diffusion model, conditioned on the text prompts $[y^{c},p_{1}^{c},\dots,p_{k}^{c}]$ and the corresponding target durations $[M,(t_{2}-t_{1}),\dots,(t_{k+1}-t_{k})]$. The denoised output associated with the base latent (guided by $y^{c}$ and $M$) serves as the final timing-controlled audio, while the $k$ denoised sub-latents (guided by $p_{j}^{c}$ and window durations) are discarded.
@@
Decoupling Attention Control modifies the query representations in text-audio cross-attention. It segments the base query latents ($q_{\text{base}}$) into window-wise queries $\{q_{j}\}_{j=1}^{k}$ according to the timing windows $[t_{j},t_{j+1}]$. To promote intra-window consistency via base-to-sub information flow, for each $j$-th sub-latent, we replace its first $(t_{j+1}-t_{j})$ positions with $q_{j}$, while keeping the other positions unchanged. Note that this targeted replacement at the sequence start aligns with the base model’s convention for handling variable-length generation.
Aggregating Attention Control enables sub-to-base information flow to enforce precise timing alignment. It operates on the attention outputs from both text-audio cross-attention and audio-only self-attention blocks. Specifically, for each sub-latent, it extracts the active segment $o_{j}^{a}$, corresponding to the initial $(t_{j+1}-t_{j})$-second positions that match the duration of the $j$-th time window. These active segments are then temporally concatenated following the sequence defined by the timing plan $[t_{j},t_{j+1}]$, resulting in a composite timing-controlled output $o_{\text{timing}}$:
@@ 4.2 Long-Form Audio Generation
Method Output Length AudioCaps MusicCaps
@@
${}^{\dagger}$ / ${}^{\dagger}$ 10 sec 1.82 1.69 - - 3.40 1.23 - 0.320
AudioLDM ${}^{\dagger}$ 10 sec 1.78 1.60 - 0.191 3.13 1.20 - 0.301
Stable ${}^{\dagger}$ 10 sec 3.60 2.32 38.27 0.340 3.51 1.32 2.93 0.480
@@ D.1 LLM Robustness on Unseen or Noisy Prompts
Timing prompts: Frying. <0.0,8.0>, Dog bakring. 0s-4s., Running water. <4.0,8.0>, Alarm ringing. From 8 to 10., Woman speaking. 8$\sim$10 sec.

### v2

@@ 3.2 Problem Formulation
We present a training-free method to adapt a pretrained DiT-based T2A model for long-form generation with timing control. We start with a base model, denoted as $f$, which takes a text prompt $y^{c}$ (i.e., overall audio caption) and a target duration $M$ as input to generate an audio waveform of $M$ seconds, where $M\in(0,M_{\text{max}})$. In practice, $M_{\text{max}}$ is typically set to 10 seconds for most existing T2A models. To enable timing control, our method incorporates an additional input: timing prompts $y^{t}=\{(p_{i}^{c},s_{i},e_{i})\}_{i=1}^{n}$. Here, each element $(p_{i}^{c},s_{i},e_{i})$ specifies that the audio content described by the fine-grained text prompt $p_{i}^{c}$ should occur within the time interval $s_{i},e_{i}]$.
Specifically, for the timing control capability, FreeAudio employs a two-stage planning-and-generation approach: 1) Planning: Leveraging LLM’s capabilities, FreeAudio first generates a detailed timing plan $y^{p}=\{(p_{j}^{c},t_{j},t_{j+1})\}_{j=1}^{k}$, using the overall text prompt $y^{c}$ and the timing prompts $y^{t}$. This plan $y^{p}$ divides the total $M$-second duration into $k$ sequential, non-overlapping windows $\{[t_{j},t_{j+1}]\}_{j=1}^{k}$ with $0=t_{1}<\cdots<t_{k+1}=M$, each associated with a recaptioned prompt $p_{j}^{c}$. 2) Generation: Based on the resulting timing plan, FreeAudio employs the Decoupling & Aggregating Attention Control to manipulate the attention outputs across the time axis, aligning each prompt $p_{j}^{c}$ with its corresponding time window $t_{j},t_{j+1}]$.
Furthermore, this planning-and-generation schema facilitates long-form audio generation, extending beyond the base model’s maximum duration. Instead of directly modeling the entire target duration $M^{\prime}>M_{\text{max}}$, which exceeds the base model’s capacity, our approach first leverages the generated timing plan $y^{p}$ to partition the task into a sequence of manageable segments, each typically having a duration less than or equal to $M_{\text{max}}$. For smoother transitions between these segments, FreeAudio introduces the Contextual Latent Composition applied at each diffusion sampling step. Additionally, FreeAudio employs the proposed Reference Guidance to facilitate the maintenance of global consistency across the full audio sequence of duration $M$’.
@@ 3.3 LLM Planning
As shown in Figure 1 (left), the LLM planning stage takes the text prompt $y^{c}$ and timing prompts $y^{t}$ to generate a detailed timing plan $y^{p}$. This plan consists of a sequence of non-overlapping time windows $t_{j},t_{j+1}]$, each associated with recaptioned prompts $p_{j}^{c}$. Specifically, we structure this planning stage using chain-of-thought (CoT) [38, 43] reasoning in two steps: timing-to-window planning to establish the time windows, followed by text recaption to assign a detailed prompt to each window.
@@
The timing-to-window planning stage is necessary for two primary reasons: 1) Temporal Overlaps: Input intervals $s_{i},e_{i}]$ in the timing prompts $y^{t}$ can overlap. Merging the text encoder’s outputs for separate overlapping prompts (e.g., $p_{x}^{c}$ and $p_{y}^{c}$) poses challenges for accurate generation guidance. In comparison, establishing non-overlapping windows shifts the fusion to the text input stage (e.g., concatenating prompts such as $p_{x}^{c}$ and $p_{y}^{c}$), which better leverages the pretrained text encoder’s capability for interpreting unified textual descriptions of combined events. 2) Temporal Gaps: The input timing prompts $y^{t}$ might not cover the entire duration, leaving gaps where no audio event is specified. By leveraging the LLM’s planning ability, this stage infers and assigns appropriate audio events for the uncovered periods within the generated timing plan $y^{p}$. This ensures temporal coherence and semantic consistency in the audio narrative, informed by the context of the planned timing windows and the overall text prompt $y^{c}$.
Specifically, the timing-to-window planning process involves several steps: First, we define the window boundaries by collecting all unique start ($s_{i}$) and end ($e_{i}$) timestamps from $y^{t}$, inserting $t_{1}=0$ and $t_{k+1}=M$, and sorting them to obtain $0=t_{1}<t_{2}<\dots<t_{k+1}=M$. This defines the non-overlapping windows $t_{j},t_{j+1}]$. Second, we initialize a preliminary description for the window-level events by aggregating all fine-grained text prompts $p_{i}^{c}$ whose original intervals $s_{i},e_{i}]$ encompass the current window $t_{j},t_{j+1}]$. Third, leveraging the overall text prompt $y^{c}$ and the current window-level event assignments, the LLM performs a planning step to systematically assign the remaining audio events from $y^{c}$ and distribute them into appropriate window slots, prioritizing the filling of uncovered intervals, thereby finalizing the overall event plan.
@@ 3.4 Timing-Controlled Generation
A naive approach to timing control might independently generate each $(t_{j+1}-t_{j})$-second audio segment using only its associated prompt $p_{j}^{c}$ from the timing plan $y^{p}$ [42]. However, simply concatenating these segments often results in poor inter-window consistency and lacks global guidance from the overall prompt $y^{c}$. To address these limitations, we construct a batch inference pipeline that jointly leverages both $y^{c}$ and $y^{p}$. Specifically, we form a batch of size $1+k$ (one base latent and $k$ sub-latents) for the base diffusion model, conditioned on the text prompts $y^{c},p_{1}^{c},\dots,p_{k}^{c}]$ and the corresponding target durations $M,(t_{2}-t_{1}),\dots,(t_{k+1}-t_{k})]$. The denoised output associated with the base latent (guided by $y^{c}$ and $M$) serves as the final timing-controlled audio, while the $k$ denoised sub-latents (guided by $p_{j}^{c}$ and window durations) are discarded.
@@
Decoupling Attention Control modifies the query representations in text-audio cross-attention. It segments the base query latents ($q_{\text{base}}$) into window-wise queries $\{q_{j}\}_{j=1}^{k}$ according to the timing windows $t_{j},t_{j+1}]$. To promote intra-window consistency via base-to-sub information flow, for each $j$-th sub-latent, we replace its first $(t_{j+1}-t_{j})$ positions with $q_{j}$, while keeping the other positions unchanged. Note that this targeted replacement at the sequence start aligns with the base model’s convention for handling variable-length generation.
Aggregating Attention Control enables sub-to-base information flow to enforce precise timing alignment. It operates on the attention outputs from both text-audio cross-attention and audio-only self-attention blocks. Specifically, for each sub-latent, it extracts the active segment $o_{j}^{a}$, corresponding to the initial $(t_{j+1}-t_{j})$-second positions that match the duration of the $j$-th time window. These active segments are then temporally concatenated following the sequence defined by the timing plan $t_{j},t_{j+1}]$, resulting in a composite timing-controlled output $o_{\text{timing}}$:
@@ 4.2 Long-Form Audio Generation
Method
Output
Length
AudioCaps MusicCaps
@@
[…]
@@ D.1 LLM Robustness on Unseen or Noisy Prompts
[…]